# Probabilistic Language-Image Pre-Training

### v2

@@ 1 Introduction
Despite their great success, most VLMs encode representations into a deterministic Euclidean space. This assumes a one-to-one correspondence between images and texts, which oversimplifies the complex nature of real-world relationships. In practice, image-text matching is inherently many-to-many. Multiple captions can accurately describe an image, each highlighting different aspects of the visual content. For example, a train image can be described by multiple captions, e.g., “a train”, “train station” or “train parked next to a station”. Conversely, a caption may correspond to several images describing similar scenes or objects, e.g., “a train” can be matched to all the train images. However, as shown in Figure 1 (b), a deterministic model (e.g., CLIP (Radford et al., 2021)) fails to capture the multiplicity, e.g., “Train Station” embedding is located to an irrelevant point to the other train images and captions. This is because the CLIP loss forces the positive pairs close and random negative pairs far away, which has no stable solution when we map them onto a point in Euclidean space.
@@
In the experiments, ProLIP slightly outperforms the deterministic CLIP model in zero-shot classification (ZSC) tasks (e.g., CLIP shows 67.2 ImageNet ZSC accuracy, while ProLIP shows 67.6). We also show the benefits of using uncertainty estimates for image-text tasks. First, we observe that our intuition and the learned uncertainty are aligned well. For example, (1) texts generally “include” images (i.e., texts are more uncertain than images), (2) shorter texts tend to be more uncertain, (3) more general texts/images tend to be more uncertain and include more specific ones (e.g., masked image and “Train Station” in Figure 1). Furthermore, we show two applications when a proper uncertainty estimate is helpful; Bayesian Prompt Re-Weighting (BPRW), a fully Bayesian approach to seek better ImageNet zero-shot prompts which improves accuracy from 74.6% 75.8%), and the uncertainty-based dataset traversal, which provides a better understanding of dataset hierarchy.
@@ 2.2 Probabilistic image-text representations
As we discussed in Section 2.1 and A.1, VL tasks also suffer from aleatoric uncertainty caused by the inherent multiplicity of image-text matching and sparse annotations. Recently, there have been attempts to tackle the inherent ambiguity of VL tasks with probabilistic embeddings (Chun et al., 2021; Ji et al., 2023; Upadhyay et al., 2023; Chun, 2024). However, these methods have a very limited scale to be used as a generic purpose VLM, such as CLIP. For example, ProbVLM (Upadhyay et al., 2023) is an ad-hoc module top on the frozen pre-trained CLIP, limiting the full exploration of the probabilistic space. Furthermore, ProbVLM is only trained on small image caption datasets, such as CUB (Wah et al., 2011) or COCO caption (Chen et al., 2015), which makes it not applicable to more practical zero-shot classification applications. MAP (Ji et al., 2023) proposes a pre-training method using a cross-attention Transformer. However, it has a limited zero-shot capability, resulting in the need to fine-tune the model for each downstream task. Furthermore, its structure is highly inefficient for retrieval systems; it needs both image and text pair to compute a similarity between them, i.e., we have to compute all possible image-text pairs to get the full similarity. Lastly, PCME++ (Chun, 2024) showed a possibility of pre-trained PrVLM, but their scalability is still limited (e.g., achieving 34% ImageNet zero-shot accuracy). We empirically observe that the objective function of PCME++ shows slow or unstable training under large-scale image-text paris. Furthermore, all these PrVLMs need heavy additional parameters to estimate uncertainty from data. ProLIP does not need a dedicated module for uncertainty estimate but employs a very efficient strategy using [UNC].
@@ 4.1 Implementation details and experimental protocol
Evaluation. We evaluate the models on 38 tasks of Datacomp evaluation suite (Gadre et al., 2024) – the full evaluation datasets are listed in Section B.3. We report five categories: ImageNet, 6 ImageNet variants with distribution shifts, 13 VTAB tasks, 3 retrieval tasks, and the average of 38 tasks. In addition, we employ the HierarCaps dataset (Alper & Averbuch-Elor, 2024) which provides captions with four different hierarchies (e.g., “water sports” “kite surfing” “kite surfer on top of the board” “kite surfer in the air on top of a red board”). Similarly, we construct new HierarImgs dataset, which provides images with four different hierarchies (See Figure B.2 for examples). We will describe the details of HierarCaps, HierarImgs, and their evaluation in Section 4.4.
@@ 4.3 Understanding the learned uncertainty
Uncertain samples visualization. From Equation 1, we can define the uncertainty of the given input by measuring , namely, (we simply denote for image uncertainty and similarly). Figure 4 shows the samples with low and high uncertainty values using this value. We extract samples from the 3.5M subset of 12.8B DataComp CommonCrawl small (Gadre et al., 2024) filtered by CLIP similarity and English filtering 111https://huggingface.co/datasets/nielsr/datacomp-small-filtered. We use ProLIP with 12.8B seen samples for the analyses.
@@ 4.4 Application using uncertainty
More examples and discussion. Due to the page limit, we include more example applications in Section C.8. Also, we discuss the limitation of diagonal covariance assumption in Appendix D.
@@ 5 Conclusion
For reproducible research, we use the open-source training dataset (DataComp 1B (Gadre et al., 2024)) and codebase (openclip (Ilharco et al., 2021)). As we clarified in Section B.3, we had 1,121,356,767 number of valid URLs among 1.39 billion URLs and 1,118,443,492 number of unique images after de-duplicating URLs (there were 147,676,246 number of duplicated URLs in the DataComp 1B URLs). All the detailed hyperparameters are clarified in Section B.1. Our results can be reproducible by our open-source implementation (https://github.com/naver-ai/prolip) and released pre-trained weights in HuggingFace (https://huggingface.co/collections/SanghyukChun/prolip-6712595dfc87fd8597350291).
Acknolwedgements
@@ A.8 Algorithm of uncertainty-aware zero-shot prompt selection
We assume a Dirichlet prior for with , where controls the “uniformity” of the posterior. If we choose large , such as 10, then will be more uniform and if we choose small , such as 0.1, then almost becomes a Direc-delta distribution. We also set the initial with the reversed uncertainty score, i.e., the normalized for each prompt. We also employ a small trick for stability. After we got the prior value, we add a small to to make the overall operation stable. Now, we explain the Expectation-Maximization (EM) algorithm for estimating the mixing proportions of a mixture of Gaussian components, incorporating a Dirichlet prior over the mixing proportions. For simplicity, we omit the class index for the remaining section.
@@ B.2 Fine-tuning details
During the project, we indeed aspire to train large models, such as ViT-H, ViT-SO400M, ViT-G or ViT-g. However, as reported by Cherti et al. (2023), ViT-H/14 CLIP model with 34B seen samples (achieving 78.0% ImageNet zero-shot accuracy) task 279 hours with 824 A100s. Using our resource (32 H100s), training the same backbone from scratch takes almost 300 days. For this reason, we tried to fine-tune the pre-trained strong backbones. Note that our goal is not to fine-tune the existing deterministic VLMs, but we report four fine-tuning results, ViT-B/16 (76.0% IN-ZS), ViT-L/16 (80.5% IN-ZS) and ViT-SO400M/14 (82.0% IN-ZS) pre-trained by SigLIP (Zhai et al., 2023) and ViT-H/14 (83.4% IN-ZS) pre-trained by DFN (Fang et al., 2024), for achieving a stronger PrVLM.
For the architectural consistency between all ProLIP models, we remove the attention pooling of pre-trained SigLIP and fine-tune the models using [CLS] token-based pooling and fix the image resolution to 224224. This will lead to slightly worse performance of the fine-tuned models than the original models. Note that we can implement both attention pooling and [UNC] token architecture simultaneously, but we did not implement the architecture for simplicity. We again emphasize that we do not aim to achieve the state-of-the-art zero-shot classification, instead, our goal is to verify whether ProLIP performs better by scaling up the base architecture even beyond ViT-B/16. For the fine-tuning, we use the learning rate as 5.0e-5, the weight decay as 0.0, and the number of seen samples as 1.28B.
@@ B.3 Training and evaluation datasets
We use 38 tasks from the DataComp evaluation suite: ImageNet (Russakovsky et al., 2015), 6 benchmarks for evaluating robustness under ImageNet distribution shifts, including ImageNet-A, ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), ImageNet v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019) and ObjectNet (Barbu et al., 2019), and 13 VTAB task (Zhai et al., 2019), including Caltech-101, CIFAR-100, CLEVR Counts, CLEVR Distance, Describable Textures, EuroSAT, KITTI Vehicle Distance, Oxford Flowers-102, Oxford-IIIT Pet, PatchCamelyon, RESISC45, SVHN and SUN397. We also employ three retrieval tasks, including Flickr (Young et al., 2014), MS-COCO Caption (Chen et al., 2015) and WinoGAViL (Bitton et al., 2022).
@@ C.5 More results for HierarImgs
Unlike text uncertainty (Figure 7), we observe that the visual uncertainty values are not discriminative by the levels in contrast to text uncertainty – See Figure 2(a). Lower-level images tend to have a larger average uncertainty (0.015) than the original images (0.014), but their differences are not significant between levels as texts. Instead of plotting every image in the same histogram, we plot the difference of uncertainty between the original image and its maksed versions in Figure C.2 (b-d).

### v4

@@ 1 Introduction
Despite their great success, most VLMs encode representations into a deterministic Euclidean space. This assumes a one-to-one correspondence between images and texts, which oversimplifies the complex nature of real-world relationships. In practice, image-text matching is inherently many-to-many (chun2025multiplicity). Multiple captions can accurately describe an image, each highlighting different aspects of the visual content. For example, a train image can be described by multiple captions, e.g., “a train”, “train station” or “train parked next to a station”. Conversely, a caption may correspond to several images describing similar scenes or objects, e.g., “a train” can be matched to all the train images. However, as shown in Figure 1 (b), a deterministic model (e.g., CLIP) fails to capture the multiplicity, e.g., “Train Station” embedding is located to an irrelevant point to the other train images and captions. This is because the CLIP loss forces the positive pairs close and random negative pairs far away, which has no stable solution when we map them onto a point in Euclidean space.
@@
In the experiments, ProLIP slightly outperforms the deterministic CLIP model in zero-shot classification (ZSC) tasks (e.g., CLIP shows 67.2 ImageNet ZSC accuracy, while ProLIP shows 67.8). We also show the benefits of using uncertainty estimates for image-text tasks. First, we observe that our intuition and the learned uncertainty are aligned well. For example, (1) texts generally “include” images (i.e., texts are more uncertain than images), (2) shorter texts tend to be more uncertain, (3) more general texts/images tend to be more uncertain and include more specific ones (e.g., masked image and “Train Station” in Figure 1). Furthermore, we show two applications when a proper uncertainty estimate is helpful; Bayesian Prompt Re-Weighting (BPRW), a fully Bayesian approach to seek better ImageNet zero-shot prompts which improves accuracy from 74.6% 75.8%), and the uncertainty-based dataset traversal, which provides a better understanding of dataset hierarchy.
@@ 2.2 Probabilistic image-text representations
As we discussed in Section 2.1 and A.1, VL tasks also suffer from aleatoric uncertainty caused by the inherent multiplicity of image-text matching and sparse annotations. Recently, there have been attempts to tackle the inherent ambiguity of VL tasks with probabilistic embeddings (Chun et al., 2021; Ji et al., 2023; Upadhyay et al., 2023; Chun, 2024). However, these methods have a very limited scale to be used as a generic purpose VLM, such as CLIP. For example, ProbVLM (Upadhyay et al., 2023) is an ad-hoc module top on the frozen pre-trained CLIP, limiting the full exploration of the probabilistic space. Furthermore, ProbVLM is only trained on small image caption datasets, such as CUB (Wah et al., 2011) or COCO caption (Chen et al., 2015), which makes it not applicable to more practical zero-shot classification applications. MAP (Ji et al., 2023) proposes a pre-training method using a cross-attention Transformer. However, it has a limited zero-shot capability, resulting in the need to fine-tune the model for each downstream task. Furthermore, its structure is highly inefficient for retrieval systems; it needs both image and text pair to compute a similarity between them, i.e., we have to compute all possible image-text pairs to get the full similarity. Lastly, PCME++ (Chun, 2024) showed a possibility of pre-trained PrVLM, but their scalability is still limited (e.g., achieving 34% ImageNet zero-shot accuracy). We empirically observe that the objective function of PCME++ shows slow or unstable training under large-scale image-text pairs. Furthermore, all these PrVLMs need heavy additional parameters to estimate uncertainty from data. ProLIP does not need a dedicated module for uncertainty estimate but employs a very efficient strategy using [UNC].
@@ 4.1 Implementation details and experimental protocol
Evaluation. We evaluate the models on 38 tasks of DataComp evaluation suite (Gadre et al., 2024) – the full benchmarks are listed in Section B.3. We report five categories: ImageNet, 6 ImageNet variants with distribution shifts, 13 VTAB tasks, 3 retrieval tasks, and the average of 38 tasks. In addition, we employ the HierarCaps dataset (Alper & Averbuch-Elor, 2024), which provides captions with four different hierarchies (e.g., “water sports” “kite surfing” “kite surfer on top of the board” “kite surfer in the air on top of a red board”). Similarly, we construct new HierarImgs dataset, which provides images with four different hierarchies (See Figure B.2 for examples). We will describe the details of HierarCaps, HierarImgs, and their evaluation in Section 4.4.
@@ 4.3 Understanding the learned uncertainty
Uncertain samples visualization. From Equation 1, we can define the uncertainty of the input by measuring , namely, (we simply denote for image uncertainty and similarly). Figure 4 shows the samples with low and high uncertainty values using this value. We extract samples from the 3.5M subset of 12.8B DataComp CommonCrawl small (Gadre et al., 2024) filtered by CLIP similarity and English filtering 111https://huggingface.co/datasets/nielsr/datacomp-small-filtered. We use ProLIP with 12.8B seen samples for the analyses.
@@ 4.4 Application using uncertainty
More examples and discussion. We include more applications in Section C.8, including filtering, dataset understanding, and LongProLIP. Appendix D discusses the diagonal covariance assumption.
@@ 5 Conclusion
For reproducible research, we use the open-source training dataset (DataComp 1B (Gadre et al., 2024)) and codebase (openclip (Ilharco et al., 2021)). As we clarified in Section B.3, we had 1,121,356,767 number of valid URLs among 1.39 billion URLs and 1,118,443,492 number of unique images after de-duplicating URLs (there were 147,676,246 number of duplicated URLs in the DataComp 1B URLs). All the detailed hyperparameters are clarified in Section B.1. Our results can be reproducible by our open-source implementation (https://github.com/naver-ai/prolip) and released pre-trained weights in HuggingFace (https://huggingface.co/collections/SanghyukChun/prolip-6712595dfc87fd8597350291); the released weights include ProLIP ViT-B/16 (from-scratch), ViT-L/16, ViT-SO400M/14, ViT-H/14 (fine-tuned), and LongProLIP ViT-B/16 fine-tuned with different datasets.
Acknowledgements
@@ A.8 Algorithm of uncertainty-aware zero-shot prompt selection
We assume a Dirichlet prior for with , where controls the “uniformity” of the posterior. If we choose large , such as 10, then will be more uniform and if we choose small , such as 0.1, then almost becomes a Dirac-delta distribution. We also set the initial with the reversed uncertainty score, i.e., the normalized for each prompt. We also employ a small trick for stability. After we got the prior value, we add a small to to make the overall operation stable. Now, we explain the Expectation-Maximization (EM) algorithm for estimating the mixing proportions of a mixture of Gaussian components, incorporating a Dirichlet prior over the mixing proportions. For simplicity, we omit the class index for the remaining section.
@@ B.2 Fine-tuning details
During the project, we indeed aspire to train large models, such as ViT-H, ViT-SO400M, ViT-G or ViT-g. However, as reported by Cherti et al. (2023), ViT-H/14 CLIP model with 34B seen samples (achieving 78.0% ImageNet zero-shot accuracy) takes 279 hours with 824 A100s. Using our resource (32 H100s), training the same backbone from scratch takes almost 300 days. For this reason, we tried to fine-tune the pre-trained strong backbones. Note that our goal is not to fine-tune the existing deterministic VLMs, but we report four fine-tuning results, ViT-B/16 (76.0% IN-ZS), ViT-L/16 (80.5% IN-ZS) and ViT-SO400M/14 (82.0% IN-ZS) pre-trained by SigLIP (Zhai et al., 2023) and ViT-H/14 (83.4% IN-ZS) pre-trained by DFN (Fang et al., 2024), for achieving a stronger PrVLM.
For the architectural consistency between all ProLIP models, we remove the attention pooling of pre-trained SigLIP and fine-tune the models using [CLS] token-based pooling and fix the image resolution to 224224. This will lead to slightly worse performance of the fine-tuned models than the original models. Note that we can implement both attention pooling and [UNC] token architecture simultaneously, but we did not implement the architecture for simplicity. We again emphasize that we do not aim to achieve the state-of-the-art performance, instead, our goal is to verify whether ProLIP performs better by scaling up the architecture beyond ViT-B/16. For the fine-tuning, we use the learning rate as 5.0e-5, the weight decay as 0.0, and the number of seen samples as 1.28B.
@@ B.3 Training and evaluation datasets
We use 38 tasks from the DataComp evaluation suite: ImageNet (Russakovsky et al., 2015), 6 benchmarks for evaluating robustness under ImageNet distribution shifts, including ImageNet-A, ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), ImageNet v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019) and ObjectNet (Barbu et al., 2019), and 13 VTAB task (Zhai et al., 2019), including Caltech-101 (Fei-Fei et al., 2004), CIFAR-100 (Krizhevsky, 2009), CLEVR Counts, CLEVR Distance (Johnson et al., 2017), Describable Textures (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), KITTI Vehicle Distance (Geiger et al., 2012), Oxford Flowers-102 (Nilsback & Zisserman, 2008), Oxford-IIIT Pet (Parkhi et al., 2012), PatchCamelyon (Veeling et al., 2018), RESISC45 (Cheng et al., 2017), SVHN (Netzer et al., 2011) and SUN397 (Xiao et al., 2010), and 3 retrieval tasks, including Flickr (Young et al., 2014), MS-COCO Caption (Chen et al., 2015) and WinoGAViL (Bitton et al., 2022). There are also 13 additional tasks, such as CIFAR-10, Country211 (Radford et al., 2021), FGVC Aircraft (Maji et al., 2013), Food-101 (Bossard et al., 2014), GTSRB (Stallkamp et al., 2012), MNIST (LeCun et al., 1998), Pascal VOC (Everingham et al., 2010), Rendered SST2 (Radford et al., 2021), STL-10 (Coates et al., 2011), iWildCam (Beery et al., 2021), FMoW (Christie et al., 2018), Dollar Street (Rojas et al., 2022), and GeoDE (Ramaswamy et al., 2024).
@@ C.5 More results for HierarImgs
Unlike text uncertainty (Figure 7), we observe that the visual uncertainty values are not discriminative by the levels in contrast to text uncertainty – See Figure 2(a). Lower-level images tend to have a larger average uncertainty (0.015) than the original images (0.014), but their differences are not significant between levels as texts. Instead of plotting every image in the same histogram, we plot the difference of uncertainty between the original image and its masked versions in Figure C.2 (b-d).
@@ C.8 More applications of the learned uncertainty
ProLIP with long context text.
Although ProLIP can capture inherent ambiguity in vision-language tasks, ProLIP is limited to capturing long context text exceeding 64 tokens. To tackle the issue, we fine-tune the pre-trained ProLIP with long context texts, following LongCLIP (Zhang et al., 2024). In our primitive study, we observe that fine-tuning solely with the long text dataset can lead to a significant performance drop for general zero-shot tasks. More detail of our extension, LongProLIP (expanding the text context length from 64 to 256), is out of the scope of this paper; we refer to a technical report for LongProLIP for more interested readers (Chun & Yun, 2025).